Circular No. 7518

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2000ek IN ESO 153-G27 R. Chassagne, Ste. Clotilde, Île de Réunion, reports his discovery of an

apparent supernova (mag 17.0) on CCD images taken with an automated 0.30-m telescope on Nov. 7.822 and 8.722 UT. SN 2000ek is located at at α = 2^{h}10^{m}30^{s}.04, δ = −53^{o}49′28′′.0 (equinox 2000.0), which is 22′′ west and 44′′ north of the center of ESO 153-G27. Nothing is visible at this location on an image taken by Chassagne on Sept. 28 (limiting mag 18), and he adds that the new object is not present on a U.K. Schmidt plate (via the Digital Sky Survey) taken on 1975 Nov. 11 (limiting mag 22).

2000 UG_{11}

M. C. Nolan, J.-L. Margot, and E. S. Howell, National Astronomy and Ionosphere Center; L. A. M. Benner, S. J. Ostro, R. F. Jurgens, and J. D. Giorgini, Jet Propulsion Laboratory; and D. B. Campbell, Cornell University, report: “Arecibo radar delay-Doppler images (2380 MHz, 12.6 cm) obtained on Nov. 3–7 show that 2000 UG_{11} is a binary system. Preliminary estimates of average diameters, based on range extents at 15-m resolution, are 230 and 100 m. The system passed through maximum range separation (i.e., projected along the line-of-sight) of ∼ 300 m within 15 min of Nov. 7.3646 UT. Goldstone radar observations (8560 MHz, 3.5 cm) on Nov. 6 confirm the binary nature of this minor planet. Using the combined data sets, we derive an orbital period of 19 ± 1.5 hr.”

PLUTO M. Sykes, University of Arizona; R. Cutri and J. Fowler, Infrared Pro-

cessing and Analysis Center, California Institute of Technology; D. Tholen, University of Hawaii; and M. Skrutskie, University of Massachusetts, report the following near-infrared photometry of Pluto from the Two-Micron All-Sky Survey facility at Cerro Tololo: 1999 Apr. 10.334 UT, J = 12.78 ± 0.03, H = 12.62 ± 0.03, K_{s} = 12.92 ± 0.03.

COMET 73P/SCHWASSMANN-WACHMANN 3 A. Nakamura, Kuma, Ehime, Japan, reports that a CCD image taken

low in the morning sky by K. Kadota (Ageo, Saitama, 0.18-m reflector) on Nov. 4.84 UT shows this comet unexpectedly bright at m_{1} = 13.2, with coma diameter 0′.5 and a 0′.8 tail in p.a. 310^{o}.

2000 November 10  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green